# Defensive Unlearning with Adversarial Training for Robust Concept Erasure in Diffusion Models

### v2

@@ 4 Effectiveness Enhancement of AdvUnlearn: Improving Tradeoff between Robustness and Utility
Warm-up: Difficulty of image generation quality retention. A straightforward implementation of AdvUnlearn (9) is to specify the upper-level optimization using ESD (3) and combine it with adversarial prompt generation (6). However, such a direct integration results in a notable decrease in image generation quality. Tab. 1 compares the performance of the vanilla ESD (i.e., concept-erased stable diffusion ) (Gandikota et al., 2023a) with its direct AT variant. The robustness of concept erasure is evaluated using ASR (attack success rate) against the adversarial prompt attack UnlearnDiffAtk (Zhang et al., 2023b). Meanwhile, the quality of image generation retention is assessed through FID. As we can see, while the direct AT variant of ESD (AT-ESD) enhances adversarial robustness with approximately a 20% reduction in ASR, it also leads to a considerable increase in FID. Fig. 3 presents visual examples of the generation produced by AT-ESD compared to the original SD v1.4 and vanilla ESD. As demonstrated, the decline in image generation authenticity under a benign prompt using AT-ESD is substantial.
Utility-retaining regularization in AdvUnlearn. We next improve the unlearning objective in AdvUnlearn (9) by explicitly prioritizing the retention of the DM’s generation utility. One potential explanation for the diminished image generation quality after AT-ESD is that ESD (3) primarily focuses on de-generating the unlearning concept in the diffusion process, thus lacking the capability to preserve image generation quality when further pressured by the robustness enhancement induced by AT. In the realm of AT for image classification, the integration of external (unlabeled) data into AT has proven to be an effective strategy for enhancing standard model utility (i.e., test accuracy) while simultaneously improving adversarial robustness. Drawing inspiration from this, we suggest the curation of a retain set comprising additional text prompts utilized for retaining model utility. Together with the ESD-based unlearning objective, we customize the upper-level optimization objective of (9) as
@@
where was defined in (5), and the second loss term penalizes the degradation of image generation quality using the current DM compared to the original under a retained concept . When selecting the retain set , it is essential to ensure that the enhancement in image generation quality does not come at the expense of the effectiveness of concept erasure, i.e., minimizing ESD loss in (10). Therefore, we utilize a large language model (LLM) as a judge to sift through these retain prompts, excluding those relevant to the targeted concept for erasure. Further details regarding the LLM judge system are available in Appx. A. We obtain these retain prompts from an external dataset, such as ImageNet (Deng et al., 2009) or COCO (Lin et al., 2014), using the prompt template ‘a photo of [OBJECT CLASS]’. As shown in Tab. 2, our proposed utility-retaining regularization effectively recovers the utility (i.e., FID of AdvUnlearn vs. that of ESD), which is otherwise compromised by AT-ESD. Yet, AdvUnlearn in Tab. 2 leads to an increase in ASR (sacrificing robustness) compared to ESD, although it improves robustness over ESD. Thus, there is room for further enhancement in AdvUnlearn. As will be shown in Sec. 5, the choice of the DM component to optimize in (9) is crucial for better performance.
@@ 5 Efficiency Enhancement of AdvUnlearn: Modularity Exploration and Fast Attack Generation
Where to robustify: Text encoder or UNet? Initially, concept erasure by ESD (5) was confined to the UNet component of a DM (Gandikota et al., 2023a). However, as shown in Tab. 2, optimizing UNet alone does not lead to sufficient robustness gain for AdvUnlearn. Moreover, there are efficiency benefits if concept erasure can be performed on the text encoder instead of UNet. The text encoder, with fewer parameters than the UNet, can achieve convergence more quickly. Most importantly, a text encoder that has undergone the unlearning process with one DM could possibly serve as a plug-in unlearner for other DMs, thereby broadening its applicability across various DMs. Furthermore, a recent study (Basu et al., 2023) demonstrates that causal components corresponding to the DM’s visual generation are concentrated in the text encoder. Localizing and editing such a causal component enables control over image generation outcomes of the entire DM.
@@
Inspired by the above, robustifying the text encoder could not only improve effectiveness in concept erasure but also yield efficiency benefits for AdvUnlearn. Tab. 3 extends Tab. 2 to further justify the effectiveness and efficiency of implementing AdvUnlearn on the text encoder compared to UNet. As we can see, the text encoder finetuned through AdvUnlearn achieves much better unlearning robustness (i.e., lower ASR) than AdvUnlearn applied to UNet (i.e., AdvUnlearn in Tab. 2), without loss of model utility as evidenced by FID. Although applying ESD to the text encoder can also improve ASR, it leads to a significant utility loss compared to its vanilla version applied to UNet (Gandikota et al., 2023a). This highlights the importance of retaining image generation quality considered in AdvUnlearn when optimizing the text encoder. In the rest of the paper, unless specified otherwise, we select the text encoder as the DM module to optimize in AdvUnlearn (9).
Fast attack generation in AdvUnlearn. Another efficiency enhancement for AdvUnlearn is to simplify the lower-level optimization of (9) using a one-step, fast attack generation method. This approach aligns with the concept of fast AT in image classification (Wong et al., 2020; Zhang et al., 2022b). The rationale is that the lower-level problem of (9) can be approximated using a quadratic program (Zhang et al., 2022b), and solving it can be achieved using the fast gradient sign method (FGSM) (Goodfellow et al., 2014; Wong et al., 2020). Specifically, let represent the perturbation added to the text prompt , e.g., via a prefix vector (Li and Liang, 2021). With an abuse of notation, we denote the perturbed prompt by , where the symbol represents the prefix attachment. FGSM determines using FGSM to solve the lower-level problem of (9):
@@
where represents random initialization, denotes the step size, and is element-wise sign operation. We refer to the utilization of one-step attack generation (11) in AdvUnlearn as its fast variant, which can also yield a substantial robustness gain in concept erasure. Tab. 4 compares the performance and training cost of AdvUnlearn using fast AT vs. (standard) AT. As we can see, the adoption of fast AT reduces the training time per iteration from to on a single NVIDIA RTX A6000 GPU, albeit with a corresponding decrease in unlearning efficacy and image generation utility. Therefore, when the need for unlearning efficacy is not exceedingly high and computational efficiency is prioritized, adopting fast AT can be an effective solution. We summarize the AdvUnlearn algorithm in Appx. B.
@@ 6.1 Experiment Setups
Concept-erasing tasks, datasets, and models. We categorize existing concept-erasing tasks (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Lyu et al., 2023; Wu et al., 2024; Wu and Harandi, 2024) into three main groups for ease of evaluation. (1) Nudity unlearning focuses on preventing DMs from generating harmful content subject to nudity-related prompts (Gandikota et al., 2023a; Zhang et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Lyu et al., 2023; Wu et al., 2024; Wu and Harandi, 2024). (2) Style unlearning aims to remove the influence of an artistic painting style in DM generation, which mimics the degeneration of copyrighted information such as the painting style (Gandikota et al., 2023a; Zhang et al., 2023a; Kumari et al., 2023a; Gandikota et al., 2023b; Lyu et al., 2023). (3) Object unlearning, akin to the previous tasks, targets the degeneration of DMs corresponding to a specific object (Gandikota et al., 2023a; Zhang et al., 2023a; Fan et al., 2023; Lyu et al., 2023; Wu et al., 2024; Wu and Harandi, 2024). The dataset for testing nudity unlearning is derived from the inappropriate image prompt (I2P) dataset (Schramowski et al., 2023), whereas the testing dataset for style unlearning is aligned with the setup described in (Gandikota et al., 2023a). In the scenario of object unlearning, GPT-4 (OpenAI, 2023) is utilized to generate distinct text prompts for each object class featured in Imagenette (Shleifer and Prokop, 2019). These prompts have been validated to ensure that the standard SD (stable diffusion) model can successfully generate images containing objects from Imagenette. Model-wise, unless specified otherwise, the pre-trained SD (Stable Diffusion) v1.4 is utilized as the base DM in concept erasing.
DM unlearning baselines. We include 8 open-sourced DM unlearning methods as our baselines: (1) ESD (erased stable diffusion) (Gandikota et al., 2023a), (2) FMN (Forget-Me-Not) (Zhang et al., 2023a), (3) AC (ablating concepts) (Kumari et al., 2023a), (4) UCE (unified concept editing) (Gandikota et al., 2023b), (5) SalUn (saliency unlearning) (Fan et al., 2023), (6) SH (ScissorHands) (Wu and Harandi, 2024), (7) ED (EraseDiff) (Wu et al., 2024), and (8) SPM (concept-SemiPermeable Membrane) (Lyu et al., 2023). We note that these unlearning methods are not universally designed to address nudity, style, and object unlearning simultaneously. Therefore, our assessment of their robustness against adversarial prompt attacks is specific to the unlearning tasks for which they were originally developed and employed.
@@
Evaluation setups. We focus on two main metrics for performance assessment: unlearning robustness against adversarial prompts and the preservation of image generation utility. For robustness evaluation, we measure the attack success rate (ASR) of DMs in the presence of adversarial prompt attacks, where a lower ASR indicates better robustness. Unless specified otherwise, we utilize UnlearnDiffAtk (Zhang et al., 2023b) for generating adversarial prompts at testing time, as it can be regarded as an unseen attack strategy different from (6) used in AdvUnlearn. Detailed settings for attack evaluation are presented in Appx. C. For utility evaluation, we use FID (Heusel et al., 2017) to assess the distributional quality of image generations. We also use CLIP score (Hessel et al., 2021) to measure their contextual alignment with prompt descriptions. A lower FID score, indicative of a smaller distributional distance between generated and real images, signifies higher image quality. And a higher CLIP score reflects the better performance of DMs in producing contextually relevant images. To compute these utility metrics, we employ DMs to generate images under prompts, randomly sampled from the COCO caption dataset (Chen et al., 2015).
@@ 6.2 Experiment Results
Robustness-utility evaluation of AdvUnlearn for nudity unlearning. In Tab. 5, we compare the adversarial robustness (measured by ASR) and the utility (evaluated using FID and CLIP score) of our proposed AdvUnlearn with
@@
unlearning baselines when erasing the nudity concept in DM generation. For ease of presentation, we also refer to the DM post-unlearning (i.e., the unlearned DM) with the name of the corresponding unlearning method. Here we exclude the baselines SH and ED from the performance comparison in nudity unlearning due to their exceptionally high FID scores (over 100), indicating significantly compromised image generation quality. Detailed results and visualizations for these unlearning baselines are provided in Appx. D. As we can see, our proposal (AdvUnlearn) demonstrates significantly improved ASR, with over 50% ASR reduction compared to ESD, except for the lowest ASR achieved by SalUn. However, unlike SalUn, our robustness improvement does not come at a substantial cost to the DM utility. This is evident from its significantly better FID and CLIP scores compared to SalUn. To complement our quantitative findings, Fig. 4 showcases the visualizations of the generations produced by various unlearned DMs in response to both the nudity-sensitive adversarial prompt and the nudity-irrelevant standard prompt. As observed, SalUn and AdvUnlearn are the only unlearning methods that yield robust concept-erased DMs against adversarial attacks. Yet, AdvUnlearn outperforms SalUn in generation quality.
@@
Effectiveness in style unlearning. In Tab. 6, we compare the robustness and utility performance of AdvUnlearn with various DM unlearning methods when removing the ‘Van Gogh’ artistic style from image generation. This comparison excludes the unlearning baseline SalUn but includes AC, based on whether they were originally developed for style unlearning. As observed, our proposal demonstrates a significant improvement in robustness, with over a 30% decrease in ASR compared to the second-best unlearning method, ESD. Crucially, this is accomplished without sacrificing model utility, as indicated by the comparable FID and CLIP scores compared to the base SD v1.4. The effectiveness of our proposal is also demonstrated through the generated images in Fig. 5 under adversarial prompt attack and ‘Van Gogh’-irrelevant benign prompt, respectively.
@@
Effectiveness in object unlearning. Tab. 7 compares the performance of AdvUnlearn with baselines when unlearning the object concept ‘Church’. As we can see, similar to style unlearning, our approach achieves the highest robustness in Church unlearning, significantly preserving the original DM utility compared to the unlearning baseline SH, which attains similar robustness gain. The superiority of AdvUnlearn can also be visualized in Fig. 6, showing DM generation examples. Other results and visualizations can be found in Appx. E.
@@
Plug-and-play capability of adversarially unlearned text encoder. Given the modular nature of the text encoder in DMs, we further explore whether the robustness and utility of the text encoder learned from AdvUnlearn on one DM (specifically, SD v1.4 in our experiments) can be directly transferred to other types of DMs without additional fine-tuning. Tab. 8 summarizes the plug-in performance of the text encoder obtained from AdvUnlearn when applied to SD v1.5, DreamShaper (Lykon, 2023a), and Protogen (Lykon, 2023b) for nudity unlearning. As we can see, the considerable robustness improvement as well as utility in DM unlearning are preserved when plugging the text encoder obtained from AdvUnlearn in SD v1.4 into other DMs (see ‘Transfer’ performance vs. ‘Original’ performance). This is most significant when transferring to SD v1.5 due to its similarity with SD v1.4. For dissimilar DMs like DreamShaper (Lykon, 2023a) and Protogen (Lykon, 2023b), the AdvUnlearn-acquired text encoder in SD v1.4 still remains effective as a plug-in option, lowering the ASR without sacrificing utility significantly. Fig. 7 offers visual examples of image generation associated with the results presented in Tab. 8.
The effect of text encoder layers on DM unlearning. In Fig. 8, we show the ASR (robustness metric) and the CLIP score (a utility metric) of post-nudity unlearning against various choices of text encoder layers for optimization in AdvUnlearn. Here the layer number equal to signifies that the first layers are optimized. We observe that the
@@
robustness gain escalates as more layers are optimized. In particular, optimizing only the initial layers failed to provide adequate robustness for DM unlearning against adversarial attacks, contrary to findings in (Basu et al., 2023), where shallower encoder layers suffice for guiding DMs in image editing albeit in a non-adversarial context. Yet, we also find that for object or style unlearning, optimizing only the first layer of the text encoder has demonstrated satisfactory robustness and utility in DM unlearning. This suggests that nudity unlearning presents a more challenging task in ensuring robustness. Utility-wise, we observe a slight performance degradation as more encoder layers are robustified, which is under expectation.
@@
Other ablation studies. In Appx. F, we demonstrate more ablation studies. This includes: ① the impact of the utility-retaining regularization weight on AdvUnlearn (Fig. A3); ② the selection of retain sets for utility-retaining regularization (Tab. A3); ③ the effect of the number of attack steps on adversarial prompt generation for AdvUnlearn (Fig. A4); ④ the impact of adversarial prompting strategy for AdvUnlearn (Tab. A4); ⑤ the robustness of SD v1.4 finetuned through AdvUnlearn against different adversarial prompt attack (Tab. A5).

### v3

@@ 4 Effectiveness Enhancement of AdvUnlearn: Improving Tradeoff between Robustness and Utility
Warm-up: Difficulty of image generation quality retention. A straightforward implementation of AdvUnlearn (9) is to specify the upper-level optimization using ESD (3) and combine it with adversarial prompt generation (6). However, such a direct integration results in a notable decrease in image generation quality. Tab. 1 compares the performance of the vanilla ESD (i.e., concept-erased stable diffusion ) [19] with its direct AT variant. The robustness of concept erasure is evaluated using ASR (attack success rate) against the adversarial prompt attack UnlearnDiffAtk [23]. Meanwhile, the quality of image generation retention
is assessed through FID. As we can see, while the direct AT variant of ESD (AT-ESD) enhances adversarial robustness with approximately a 20% reduction in ASR, it also leads to a considerable increase in FID. Fig. 3 presents visual examples of the generation produced by AT-ESD compared to the original SD v1.4 and vanilla ESD. As demonstrated, the decline in image generation authenticity under a benign prompt using AT-ESD is substantial.
Utility-retaining regularization in AdvUnlearn. We next improve the unlearning objective in AdvUnlearn (9) by explicitly prioritizing the retention of the DM’s generation utility. One potential explanation for the diminished image generation quality after AT-ESD is that ESD (3) primarily focuses on de-generating the unlearning concept in the diffusion process, thus lacking the capability to preserve image generation quality when further pressured by the robustness enhancement induced by AT. In the realm of AT for image classification, the integration of external (unlabeled) data into AT has proven to be an effective strategy for enhancing standard model utility (i.e., test accuracy) while simultaneously improving adversarial robustness [32]. Drawing inspiration from this, we suggest the curation of a retain set comprising additional text prompts utilized for retaining model utility. Together with the ESD-based unlearning objective, we customize the upper-level optimization objective of (9) as
where was defined in (5), and the second loss term penalizes the degradation of image generation quality using the current DM compared to the original under a retained concept .
When selecting the retain set , it is essential to ensure that the enhancement in image generation quality does not come at the expense of the effectiveness of concept erasure, i.e., minimizing ESD loss in (10). Therefore, we utilize a large language model (LLM) as a judge to sift through these retain prompts, excluding those relevant to the targeted concept for erasure. Further details regarding the LLM judge system are available in Appx. A. We obtain these retain prompts from an external dataset, such as ImageNet [100] or COCO [101], using the prompt template ‘a photo of [OBJECT CLASS]’. The finalized retain set consists of distinct prompts. During training, a prompt batch of size randomly selected from in support of utility-retaining regularization. The primary goal of is not to train the model on producing specific objects or concepts; Instead, it aims to guide the model in generating general, non-forgetting content effectively. As will be evidenced in Figs. 4-6, incorporating enhances the general utility of the unlearned DM during the testing phase. Test-time prompts in these figures include varied objects like ‘toilet’, ‘Picasso’, and ‘cassette player’ not part of , demonstrating the unlearned model’s generalization capabilities.
@@
As shown in Tab. 2, our proposed utility-retaining regularization effectively recovers the utility (i.e., FID of AdvUnlearn vs. that of ESD), which is otherwise compromised by AT-ESD. Yet, AdvUnlearn in Tab. 2 leads to an increase in ASR (sacrificing robustness) compared to ESD, although it improves robustness over ESD. Thus, there is room for further enhancement in AdvUnlearn. As will be shown in Sec. 5, the choice of the DM component to optimize in (9) is crucial for better performance.
@@ 5 Efficiency Enhancement of AdvUnlearn: Modularity Exploration and Fast Attack Generation
Where to robustify: Text encoder or UNet? Initially, concept erasure by ESD (5) was confined to the UNet component of a DM [19]. However, as shown in Tab. 2, optimizing UNet alone does not lead to sufficient robustness gain for AdvUnlearn. Moreover, there are efficiency benefits if concept erasure can be performed on the text encoder instead of UNet. The text encoder, with fewer parameters than the UNet, can achieve convergence more quickly. Most importantly, a text encoder that has undergone the unlearning process with one DM could possibly serve as a plug-in unlearner for other DMs, thereby broadening its applicability across various DMs. Furthermore, a recent study [102] demonstrates that causal components corresponding to the DM’s visual generation are concentrated in the text encoder. Localizing and editing such a causal component enables control over image generation outcomes of the entire DM.
@@
Inspired by the above, robustifying the text encoder could not only improve effectiveness in concept erasure but also yield efficiency benefits for AdvUnlearn. Tab. 3 extends Tab. 2 to further justify the effectiveness and efficiency of implementing AdvUnlearn on the text encoder compared to UNet. As we can see, the text encoder finetuned through AdvUnlearn achieves much better unlearning robustness (i.e., lower ASR) than AdvUnlearn applied to UNet (i.e., AdvUnlearn in Tab. 2), without loss of model utility as evidenced by FID. Although applying ESD to the text encoder can also improve ASR, it leads to a significant utility loss compared to its vanilla version applied to UNet [19]. This highlights the importance of retaining image generation quality considered in AdvUnlearn when optimizing the text encoder. In the rest of the paper, unless specified otherwise, we select the text encoder as the DM module to optimize in AdvUnlearn (9).
Fast attack generation in AdvUnlearn. Another efficiency enhancement for AdvUnlearn is to simplify the lower-level optimization of (9) using a one-step, fast attack generation method. This approach aligns with the concept of fast AT in image classification [87, 92]. The rationale is that the lower-level problem of (9) can be approximated using a quadratic program [92], and solving it can be achieved using the fast gradient sign method (FGSM) [71, 87]. Specifically, let represent the perturbation added to the text prompt , e.g., via a prefix vector [103]. With an abuse of notation, we denote the perturbed prompt by , where the symbol represents the prefix attachment. FGSM determines using FGSM to solve the lower-level problem of (9):
@@
where represents random initialization, denotes the step size, and is element-wise sign operation. We refer to the utilization of one-step attack generation (11) in AdvUnlearn as its fast variant, which can also yield a substantial robustness gain in concept erasure. Tab. 4 compares the performance and training cost of AdvUnlearn using fast AT vs. (standard) AT, where the attack step of standard AT is set to 30. As we can see, the adoption of fast AT reduces the training time per iteration from to on a single NVIDIA RTX A6000 GPU, albeit with a corresponding decrease in unlearning efficacy and image generation utility. Therefore, when the need for unlearning efficacy is not exceedingly high and computational efficiency is prioritized, adopting fast AT can be an effective solution. We summarize the AdvUnlearn algorithm in Appx. B.
@@ 6.1 Experiment Setups
Concept-erasing tasks, datasets, and models. We categorize existing concept-erasing tasks [19, 20, 21, 22, 27, 28, 53, 54] into three main groups for ease of evaluation. (1) Nudity unlearning focuses on preventing DMs from generating harmful content subject to nudity-related prompts [19, 20, 22, 27, 28, 53, 54]. (2) Style unlearning aims to remove the influence of an artistic painting style in DM generation, which mimics the degeneration of copyrighted information such as the painting style [19, 20, 21, 22, 28]. (3) Object unlearning, akin to the previous tasks, targets the degeneration of DMs corresponding to a specific object [19, 20, 27, 28, 53, 54]. The dataset for testing nudity unlearning is derived from the inappropriate image prompt (I2P) dataset [10], whereas the testing dataset for style unlearning is aligned with the setup described in [19]. In the scenario of object unlearning, GPT-4 [104] is utilized to generate distinct text prompts for each object class featured in Imagenette [105]. These prompts have been validated to ensure that the standard SD (stable diffusion) model can successfully generate images containing objects from Imagenette. Model-wise, unless specified otherwise, the pre-trained SD (Stable Diffusion) v1.4 is utilized as the base DM in concept erasing.
DM unlearning baselines. We include 8 open-sourced DM unlearning methods as our baselines: (1) ESD (erased stable diffusion) [19], (2) FMN (Forget-Me-Not) [20], (3) AC (ablating concepts) [21], (4) UCE (unified concept editing) [22], (5) SalUn (saliency unlearning) [27], (6) SH (ScissorHands) [54], (7) ED (EraseDiff) [53], and (8) SPM (concept-SemiPermeable Membrane) [28]. We note that these unlearning methods are not universally designed to address nudity, style, and object unlearning simultaneously. Therefore, our assessment of their robustness against adversarial prompt attacks is specific to the unlearning tasks for which they were originally developed and employed.
@@
Evaluation setups. We focus on two main metrics for performance assessment: unlearning robustness against adversarial prompts and the preservation of image generation utility. For robustness evaluation, we measure the attack success rate (ASR) of DMs in the presence of adversarial prompt attacks, where a lower ASR indicates better robustness. Unless specified otherwise, we utilize UnlearnDiffAtk [23] for generating adversarial prompts at testing time, as it can be regarded as an unseen attack strategy different from (6) used in AdvUnlearn. Detailed settings for attack evaluation are presented in Appx. C. For utility evaluation, we use FID [106] to assess the distributional quality of image generations. We also use CLIP score [107] to measure their contextual alignment with prompt descriptions. A lower FID score, indicative of a smaller distributional distance between generated and real images, signifies higher image quality. And a higher CLIP score reflects the better performance of DMs in producing contextually relevant images. To compute these utility metrics, we employ DMs to generate images under prompts, randomly sampled from the COCO caption dataset [108].
@@ 6.2 Experiment Results
Robustness-utility evaluation of AdvUnlearn for nudity unlearning. In Tab. 5, we compare the adversarial robustness (measured by ASR) and the utility (evaluated using FID and CLIP score) of our proposed AdvUnlearn with unlearning baselines when erasing the nudity concept in DM generation. For ease of presentation, we also refer to the DM post-unlearning (i.e., the unlearned DM) with the name of the corresponding unlearning method. Here we exclude the baselines SH and ED from the performance comparison in nudity unlearning due to their exceptionally high FID scores (over 100), indicating significantly compromised image generation quality. Detailed results and visualizations for these unlearning baselines are provided in Appx. D. As we can see, our proposal (AdvUnlearn)
@@
demonstrates significantly improved ASR, with over 50% ASR reduction compared to ESD, except for the lowest ASR achieved by SalUn. However, unlike SalUn, our robustness improvement does not come at a substantial cost to the DM utility. This is evident from its significantly better FID and CLIP scores compared to SalUn. To complement our quantitative findings, Fig. 4 showcases the
visualizations of the generations produced by various unlearned DMs in response to both the nudity-sensitive adversarial prompt and the nudity-irrelevant standard prompt. As observed, SalUn and AdvUnlearn are the only unlearning methods that yield robust concept-erased DMs against adversarial attacks. Yet, AdvUnlearn outperforms SalUn in generation quality.
@@
Effectiveness in style unlearning. In Tab. 6, we compare the robustness and utility performance of AdvUnlearn with various DM unlearning methods when removing the ‘Van Gogh’ artistic style from image generation. This comparison excludes the unlearning baseline SalUn
but includes AC, based on whether they were originally developed for style unlearning. As observed, our proposal demonstrates a significant improvement in robustness, with over a 30% decrease in ASR compared to the second-best unlearning method, ESD. Crucially, this is accomplished without sacrificing model utility, as indicated by the comparable FID and CLIP scores compared to the base SD v1.4. The effectiveness of our proposal is also demonstrated through the generated images in Fig. 5 under adversarial prompt attack and ‘Van Gogh’-irrelevant benign prompt, respectively.
@@
Effectiveness in object unlearning. Tab. 7 compares the performance of AdvUnlearn with baselines when unlearning the object concept ‘Church’. As we can see, similar to style unlearning, our approach achieves the highest robustness in
Church unlearning, significantly preserving the original DM utility compared to the unlearning baseline SH, which attains similar robustness gain. The superiority of AdvUnlearn can also be visualized in Fig. 6, showing DM generation examples. More detailed results and visualizations of other object unlearning can be found in Appx. E.
@@
Plug-and-play capability of adversarially unlearned text encoder. Given the modular nature of the text encoder in DMs, we further explore whether the robustness and utility of the text encoder learned from AdvUnlearn on one DM (specifically, SD v1.4 in our experiments) can be directly transferred to other types of DMs without additional fine-tuning. Tab. 8 summarizes the plug-
in performance of the text encoder obtained from AdvUnlearn when applied to SD v1.5, DreamShaper [109], and Protogen [110] for nudity unlearning. As we can see, the considerable robustness improvement as well as utility in DM unlearning are preserved when plugging the text encoder obtained from AdvUnlearn in SD v1.4 into other DMs (see ‘Transfer’ performance vs. ‘Original’ performance). This is most significant when transferring to SD v1.5 due to its similarity with SD v1.4. For dissimilar DMs like DreamShaper [109] and Protogen [110], the AdvUnlearn-acquired text encoder in SD v1.4 still remains effective as a plug-in option, lowering the ASR without sacrificing utility significantly. Fig. 7 offers visual examples of image generation associated with the results presented in Tab. 8.
@@
The effect of text encoder layers on DM unlearning. In Fig. 8, we show the ASR (robustness metric) and the CLIP score (a utility metric) of post-nudity unlearning against various choices of text encoder layers for optimization in AdvUnlearn. Here the layer number equal to signifies that the first layers are optimized. We observe that the robustness gain escalates as more layers are optimized. In particular, optimizing only the initial layers failed to provide adequate robustness for DM unlearning against adversarial attacks, contrary to findings in [102], where shallower encoder layers suffice for guiding DMs in image editing albeit in a non-adversarial context. Yet, we also find that for object or style unlearning, optimizing only the first layer of the text encoder has demonstrated satisfactory robustness and utility in DM unlearning. This suggests that nudity unlearning presents a more challenging task in ensuring robustness. Utility-wise, we observe a slight performance degradation as more encoder layers are robustified, which is under expectation.
| Attack Method | Nudity | Van Gogh | Church | Parachute | Tench | Garbage Truck |
| UnlearnDiffAtk | 21.13% | 2% | 6% | 14% | 4% | 8% |
| CCE | 39.44% | 28% | 36% | 48% | 24% | 44% |
| PEZ | 3.52% | 0% | 2% | 0% | 0% | 4% |
| PH2P | 5.63% | 0% | 4% | 0% | 2% | 6% |
Choice of adversarial attacks. In Tab. 9, we show the attack success rate (ASR), the robustness metric of post-nudity unlearning against various choices of adversarial prompt attacks: ① CCE (circumventing concept erasure) [111] utilizes textual inversion to generate universal adversarial attacks in the embedding space. By inverting an erased concept into a ‘new’ token embedding, learned from multiple images featuring the target concept, this embedding is then inserted into the target text prompt. ② PEZ (hard prompts made easy) [111] is to generate an entire text prompt for the target image by optimizing through the cosine similarity. ③ PH2P (prompting hard or hardly prompting) [112] is similar to PEZ but with different optimization objective. ④ UnlearnDiffAtk [23] has been used as the default method for generating attacks in this work. When the attack is based on discrete prompts (such as UnlearnDiffAtk, PEZ, and PH2P), our proposed method AdvUnlearn consistently achieves remarkable erasure performance and robustness. Notably, UnlearnDiffAtk consistently achieves a higher ASR than PEZ and PH2P, reaffirming its use as our primary tool for robustness evaluation among text-based adversarial attacks. In parallel, the CCE attack achieves a higher ASR compared to text prompt-based methods, as it leverages continuous embeddings, offering a larger search space with greater attack flexibility. This is not surprising as the textual inversion is engineered to learn a ‘new’ continuous token embedding, enabling the representation of objects not encountered during training.
@@
Other ablation studies. In Appx. F, we demonstrate more ablation studies. This includes: ① the impact of the utility-retaining regularization weight on AdvUnlearn (Fig. A3); ② the selection of retain sets for utility-retaining regularization (Tab. A3); ③ the impact of adversarial prompting strategy for AdvUnlearn (Tab. A4); ④ the robustness of SD v1.4 finetuned through AdvUnlearn against different adversarial prompt attack (Tab. A5).